OpenStudy (falling_in_katt):

Simplify the expression where possible. \[(X ^{3}Y ^{2})^{4}\]

OpenStudy (condor34):

4x^3y^2

OpenStudy (campbell_st):

the index law for power of a power is multiply the powers \[(x^a)^b = x^{a \times b}\] so the power of 4 outside the brackets operates on both terms inside \[x^{3 \times 4}y^{2 \times 4}\] just simplify it

OpenStudy (falling_in_katt):

\[x ^{12}y ^{8}?\]

OpenStudy (campbell_st):

that's correct